# Inference-Aware Prompt Optimization for Aligning Black-Box Large Language Models

### v1

@@ Introduction
16:  $\mathcal{D}\leftarrow\textsc{Batch-Query}\bigl{(}\mathcal{B})$
17:  $Q^{\alpha}_{(r)}\leftarrow\textsc{Estimate-Q}(\mathcal{D})$
18:  for all $c\in\mathcal{C}$ do
19:   $\mathcal{A}_{c}^{(r)}\leftarrow\{a:\text{F}_{c,a}=\textbf{true}\}$
20:   Rank $\mathcal{A}_{c}^{(r)}$ by $Q^{\alpha}_{(r)}(c,a)$
21:   Remove bottom $\lceil|\mathcal{A}_{c}^{(r)}|/2\rceil$ arms // i.e. update $\mathbf{F}$
22:  end for
@@
24:return $\{a_{c}^{\star}\}_{c\in\mathcal{C}}$ // one survivor per context
@@
Algorithm 1 proceeds in $R=\lceil\log_{2}|\mathcal{A}|\rceil$ rounds, and tracks per context active arm using the flag $\mathbf{F}$. Each round is allocated an equal pull budget of $n_{r}=\lfloor T/R\rfloor$. An allocation routine, Allocate(F, $n_{r}$), divides this budget among the current set of unique active arms, aggregated across all contexts. Based on this allocation, a batch of inference calls is issued to the target LLM. The resulting completions are scored using a reward function or verifier and stored in the dataset $\mathcal{D}$. The $Q$-values are then estimated from the collected data. Within each context, arms are ranked and the worst performing half are eliminated. After all rounds are completed, the algorithm returns a single final arm for each context.
@@
$\lambda(a)=\begin{cases}(\lfloor\frac{n_{r}N^{(r)}_{p,max}}{M}\rfloor&\text{if }a=(p,N_{p}^{\max})\in\mathcal{A}^{(r)},\\ 0&\text{otherwise},\end{cases}$ (6)
where $M=\sum_{p:(p,N_{p}^{\max})\in\mathcal{A}^{(r)}}N^{(r)}_{p,\max}$ is the total cost of sampling all such maximal arms once. This policy maintains uniform coverage over prompts while respecting cost asymmetries and ensures that the maximum scale of every prompt has an equal number of samples.
We now derive error bounds [^1] Proof in the appendix for Psst under the allocation policies described above.
@@
Let $R=\lceil\log_{2}|\mathcal{A}|\rceil$ be the number of trimming rounds, and $o_{k}^{\min},o_{k}^{\max}]=[-1,1]$ and define the cost–gap complexity
$H_{1}^{c}=\max_{(c,a_{i})\neq a^{c}_{1}}\frac{\bar{N}_{max}}{\Delta_{c,a^{i}}^{\,2}},\quad H_{1}=\max_{c}H_{1}^{c}.$
$H_{2}^{c}=\max_{(c,a_{i})\neq a^{c}_{1}}\frac{i\bar{N}_{max}}{\Delta_{c,a^{i}}^{\,2}},\quad H_{2}=\max_{c}H_{2}^{c}.$
$\text{where},\ \Delta_{c,a^{i}}=Q^{\alpha}_{c,a^{1}}-Q^{\alpha}_{c,a^{i}}$, arms are indexed based on ascending order of $Q^{\alpha}_{c,a}$ under that context and $\bar{N}_{max}=\frac{a_{1}(N)+N_{max}}{2}$. Running PSST with the structure-aware allocation of for a total prompt complication $T$ returns the optimal arm in every context with probability at least
$1-3|\mathcal{C}|R\exp\!\Big{(}-\tfrac{T}{min(2|\mathcal{P}|H_{1},8|C|H_{2})R}\Big{)}.$
@@
$T=O\Big{(}min(|\mathcal{P}|H_{1},|\mathcal{C}|H_{2})R\,\log\!\Bigl{(}\tfrac{|\mathcal{C}|R}{\delta}\Bigr{)}\Big{)}.$
Note that applying Sequential-Halving without leveraging the structure of Iapo—specifically, without any form of information sharing across scales or contexts—incurs a sample complexity of $O(|\mathcal{C}|N_{\max})$ higher.
@@
While we describe the algorithm as where we use a new set of data in each round $\mathcal{D}$, it has been shown that in similar halving-style algorithms (Fabiano and Cazenave 2021), data accumulating all past observations—known as stockpiling—can improve the complexity of $T$ by reducing the outer $R$-factor, and is recommended to use with Psst.
@@
To further reduce the budget requirement of Psst, we introduce Top-$\boldsymbol{K}$ Screening, a practical heuristic that executes a short, uniform prompt screening at unit scale to trim obviously suboptimal prompts before running full PSST. Top-$\boldsymbol{K}$ Screening takes a budget fraction $T_{0}=\lfloor\rho T\rfloor$ ($\rho\in(0,1)$) from Psst. With scale restriction of N=1N{=}1, the budget is allocated uniformly across prompts: each $p\in\mathcal{P}$ receives $\big{\lfloor}T_{0}/|\mathcal{P}|\big{\rfloor}$ i.i.d. samples. Based on this data, $Q^{\alpha}(c,p,1)$ is estimated $\forall c\in\mathcal{C},p\in\mathcal{P}$. For each context $c$, we retain the $K$ best prompts $\mathcal{P}^{(0)}_{c}=\mathrm{Top}\text{-}K\{\,\widehat{Q}^{\alpha}(c,p,1):p\in\mathcal{P}\,\}$ and discard the rest. The subsequent PSST run is then restricted to the reduced arm sets $\mathcal{A}^{(1)}_{c}=\{(p,N):p\in\mathcal{P}^{(0)}_{c},\ N\in\mathcal{N}\}$ for each $c$, and uses the remaining budget $T^{\prime}=T-T_{0}$. In the next section, we demonstrate that the screening strategy can significantly improve performance in low training budget settings without compromising quality for practical tasks. However, theoretical guarantees comparable to those of full Psst cannot be established; counterexample tasks can be carefully constructed within Iapo framework, where Top-K screening will behave suboptimally for any $K<|\mathcal{P}|$.
@@
  • To demonstrate that Psst and the Top-K Screening heuristic are highly effective at learning policy $\pi_{T}$.
  • To show that Iapo improves the average cost-adjusted reward (ACR) compared to inference strategy agnostic optimization.
@@
We compare Psst and Top-K Screening with several baselines. We denote Top-K Screening with $K=1,K=4$,and $K=8$ as Psst+K1+K1, Psst+K4+K4, and Psst+K8+K8 respectively. For these heuristics, we fix $\rho=0.2$, which was found to perform best across all datasets. Full Psst is parameter-free and does not require any tuning. In our first set of experiments, we compare our proposed methods against several standard exploration strategies:
@@
  • UCB: At each turn, selects the arm with the highest optimistic $Q$ estimate. The exploration constant $0.1$ after tuning.
@@
Note that all hyperparameter sweep results are in the supplementary material; we report results with the best setting found across all six datasets.
@@
Environments $\alpha$ $\lvert\mathcal{P}\rvert$ $N_{\max}$ $o_{k}^{max}$ $|\mathcal{X}|$ $|\mathcal{C}|$
@@
We evaluated inference-aware optimization across a total of six environments. Key details are provided in Table 1. Environments one and four are synthetically constructed to mimic Iapo tasks, where prompt-query pair score distributions $o_{i}(c,P,1)$ are modeled using categorical distributions. We introduce them to validate some of the theoretical findings. The remaining four environments are based on widely-used real-world datasets. Among these, MATH (Hendrycks et al. 2021) and CommonsenseQA (Talmor et al. 2018) are used to evaluate reasoning tasks under Majority Voting (MV), while Helpful-Harmless (Bai et al. 2022) and Summarization (Stiennon et al. 2020) are chosen for Best-of-N (BoN) evaluation.
For the MV tasks, the task objective is defined as an exact match with the correct answer. All three BoN tasks are bi-objective, and we use publicly available reward models from previous multi-objective LLM alignment studies to score completions (see appendix for links). The cost objective in all six tasks is defined to be proportional to the average number of tokens per response. For context specification, MV tasks include a budget regime \$low, mid, high\$, while BoN tasks include both the budget and the bi-objective weights, which range from 0.1 to 0.9 for each objective. For example, in the helpful–harmless task, a context might be represented as $\{\text{helpful}:0.3,\text{harmless}:0.7,\text{budget}:\text{high}(1.0)\}$. Further details, including all prompts, are provided in the supplementary material.
To construct the environments, we first generated a set of instruction prompts for each task using ChatGPT-O3. We then generated 128 responses for each prompt–query pair and estimated the score distribution using a categorical model. All completions were produced using the LLaMA-3.3-70B-Instruct, a widely used open-source model (Meta_AI 2024), which we treat as a black-box throughout our experiments. Generation was carried out with vLLM(Kwon et al. 2023) on a cluster of 8 A100 GPUs, totaling approximately 2,000 GPU-hours. Once the environments are constructed, all experiments can be run via a standard CPU quickly. We will publish the environments and code with the paper, enabling full reproducibility without any substantial computational resources.
@@
All reported curves are averages over 200 independent runs. For synthetic environments, we instantiate 200 independent environments and report the average performance across them. For the remaining four environments, each run reshuffles the dataset, performs an $80/20$ train–test split, and trains the policy on the training set. In all six environments, we evaluate ACR on the test set using $10,000$ samples. Performance for each budget is the mean across the 200 runs, with standard error of the mean (SEM) error bars. Statistical significance is assessed using the Wilcoxon paired two-sided test with alpha $0.05$, and we indicate when differences are significant in the discussion. The full set of results is in the appendix.
Comparison of Exploration Strategies (Fig. 3).
@@
Psst and the Top-K screening heuristic consistently outperform all baselines. Across all six domains, where the per-context action spaces are large ($|\mathcal{P}|N_{\max}\in[640,1536]$), UCB, softmax, and $\varepsilon$-greedy methods struggle to explore effectively. Among the baselines, UCB performs comparably in some domains after $T=20\text{K}$, but only with extensive hyperparameter tuning. Furthermore, these baselines are fully sequential and cannot leverage the cost and computational efficiency benefits of batch exploration. Full Psst attains the best final performance across four settings, while Psst+KXK\!X typically reaches strong policies faster, matching or exceeding Psst on three of the four real‑data tasks when the budget is small. Under aggressive pruning (small $K$), however, the heuristic becomes suboptimal—most notably on summarization and on the synthetic benchmarks—suggesting that Psst+KXK\!X is attractive under tight budgets, whereas full Psst is preferable for critical tasks such as long‑horizon, high‑frequency deployment. Finally, the statistical test also validates that Psst, along with Top-K screening, significantly outperforms baselines in all six datasets and under nearly all budgets. These findings indicate that our approach reliably discovers well-aligned solutions using as few as $5\text{K}$ to $20\text{K}$ inference calls in practical settings.
Importance of Inference-Awareness (Fig. 4).
@@
We examine the role of inference awareness in prompt optimization. Across all six datasets, Iapo methods markedly outperform the inference-agnostic methods, demonstrating the gains achievable when jointly optimizing the prompt and inference scale. $\mathit{TRIPLE}(N=1)$ fails as it does not leverage inference scaling. On the other hand, $\mathit{TRIPLE}(N=Random)$ fails because it does not optimize the scaling for different contexts. The screening variant PSST+K1+K1—which effectively approximates a near‑decoupled (prompt‑only) procedure—fails to reach the optimum in most cases, performing competitively only on CommonsenseQA and showing pronounced underperformance on summarization. This is because it gets stuck with deceiving prompts that fail to scale compared to prompts that may not perform well under single-shot but improve significantly under scaling. These findings underscore the essential role of Iapo in aligning black‑box LLMs and the pitfalls of disjoint optimization. Overall, Iapo outperforms disjoint optimization by up to 25% and prompt-only optimization by up to 50%.
@@
We present an inference-aware prompt optimization (Iapo) framework for aligning black-box LLMs, emphasizing that prompts and deployment-time inference scaling strategy are tightly coupled and should be optimized jointly. Our proposed Psst and Top-K Screening heuristic demonstrate consistent improvements over strong baselines across six different settings. Looking ahead, we plan to explore richer inference-scaling policies (e.g., adaptive BoN/MV schedules, stopping rules, and tree search). We also aim to extend the framework to multi-objective alignment with explicit cost/latency constraints and to study long-horizon deployments under distribution shift.
@@ Acknowledgments
Let $R=\lceil\log_{2}|\mathcal{A}|\rceil$ be the number of trimming rounds, and $o_{k}^{\min},o_{k}^{\max}]=[-1,1]$ and define the cost–gap complexity
$H_{1}^{c}=\max_{(c,a_{i})\neq a^{c}_{1}}\frac{\bar{N}_{max}}{\Delta_{c,a^{i}}^{\,2}},\quad H_{1}=\max_{c}H_{1}^{c}.$
$H_{2}^{c}=\max_{(c,a_{i})\neq a^{c}_{1}}\frac{i\bar{N}_{max}}{\Delta_{c,a^{i}}^{\,2}},\quad H_{2}=\max_{c}H_{2}^{c}.$
$\text{where},\ \Delta_{c,a^{i}}=Q^{\alpha}_{c,a^{1}}-Q^{\alpha}_{c,a^{i}}$, arms are indexed based on ascending order of $Q^{\alpha}_{c,a}$ under that context and $\bar{N}_{max}=\frac{a_{1}(N)+N_{max}}{2}$. Running PSST with the structure-aware allocation of for a total prompt complication $T$ returns the optimal arm in every context with probability at least
$1-3|\mathcal{C}|R\exp\!\Big{(}-\tfrac{T}{min(2|\mathcal{P}|H_{1},8|C|H_{2})R}\Big{)}.$
@@
$T=O\Big{(}min(|\mathcal{P}|H_{1},|\mathcal{C}|H_{2})R\,\log\!\Bigl{(}\tfrac{|\mathcal{C}|R}{\delta}\Bigr{)}\Big{)}.$
@@
Assume that the best arm was not eliminated before round $r$. Then due to Hoeffding’s inequality for any arm $a^{i}\in\mathcal{A}^{(r)}_{c}$,
$\Pr\!\big{[}\hat{Q}^{\alpha,(r)}_{c,a_{1}}<\hat{Q}^{\alpha,(r)}_{c,a_{i}}\big{]}\;\leq\;\exp\!\left(-\frac{1}{2}harmonic(t_{r_{1}},t_{r_{i}})\,\Delta_{c,a_{i}}^{2}\right).$
Here, $t_{r}$ is the number of samples that were used to estimate the $Q$ value. Letting $N_{r}$ denote the number of arms in $A^{(r)}_{c}$ whose empirical average is larger than that of the optimal arm, we have:
$\mathbb{E}[N_{r}]$ $=\sum_{a_{i}\in\mathcal{A}^{(r)}_{c}}\Pr\!\big{[}\hat{Q}^{\alpha,(r)}_{c,a_{1}}<\hat{Q}^{\alpha,(r)}_{c,a_{i}}\big{]}$
$\leq\sum_{a_{i}\in\mathcal{A}^{(r)}_{c}}\exp\!\Big{(}-\frac{1}{2}harmonic(t_{r_{1}},t_{r_{i}})\Delta_{c,a_{i}}^{2}\Big{)}$
$\leq\sum_{a_{i}\in\mathcal{A}^{(r)}_{c}}\exp\!\Big{(}-\Delta_{c,a_{i}}^{2}\cdot\tfrac{T}{2|\mathcal{P}|\bar{N}_{i}\log_{2}|\mathcal{A}|}\Big{)}$
$\leq\;|\mathcal{A}^{(r)}_{c}|\max_{i\in\mathcal{A}^{(r)}_{c}}\exp\!\Big{(}-\Delta_{c,a_{i}}^{2}\cdot\tfrac{T}{2|\mathcal{P}|\bar{N}_{max}\log_{2}|\mathcal{A}|}\Big{)}$
$\leq\;|\mathcal{A}^{(r)}_{c}|\exp\!\Big{(}-\tfrac{T}{2|\mathcal{P}|H_{1}^{c}R}\Big{)}$
[…]
$\Pr\!\Big{[}N_{r}>\tfrac{1}{2}|\mathcal{A}^{(r)}_{c}|\Big{]}\;\leq\;2\,\mathbb{E}[N_{r}]/|\mathcal{A}^{(r)}_{c}|\;\leq\;2\exp\!\Big{(}-\tfrac{T}{2|\mathcal{P}|H_{1}^{c}R}\Big{)}$
and the lemma follows. ∎

### v2

@@ Introduction
16:  end for
17:  $\mathcal{D}\leftarrow\textsc{Batch-Query}\bigl(\mathcal{B})$
18:  $Q^{\alpha}_{(r)}\leftarrow\textsc{Estimate-Q}(\mathcal{D})$
19:  for all $c\in\mathcal{C}$ do
20:   $\mathcal{A}_{c}^{(r)}\leftarrow\{a:\text{F}_{c,a}=\textbf{true}\}$
21:   Rank $\mathcal{A}_{c}^{(r)}$ by $Q^{\alpha}_{(r)}(c,a)$
22:   Remove bottom $\lceil|\mathcal{A}_{c}^{(r)}|/2\rceil$ arms // i.e. update $\mathbf{F}$
@@
24:end for
25:return $\{a_{c}^{\star}\}_{c\in\mathcal{C}}$ // one survivor per context
@@
Algorithm 1 proceeds in $R=\lceil\log_{2}|\mathcal{A}|\rceil$ rounds, and tracks per context active arm using the flag $\mathbf{F}$. Each round is allocated an equal pull budget of $n_{r}=\lfloor T/R\rfloor$. An allocation routine, Allocate(F, $n_{r}$), divides this budget among the current set of unique active arms, aggregated across all contexts. Based on this allocation, a batch of inference calls is issued to the target LLM. The resulting completions are scored using a reward function or verifier and stored in the dataset $\mathcal{D}$. The $Q$-values are then estimated from the collected data. Within each context, arms are ranked, and the worst-performing half are eliminated. After all rounds are completed, the algorithm returns a single final arm for each context.
@@
$\lambda^{(r)}(a)=\begin{cases}\lfloor\frac{n_{r}}{M}\rfloor&\text{if }a=(p,N^{(r)}_{p,\max})\in\mathcal{A}^{(r)},\\ 0&\text{otherwise},\end{cases}$ (6)
where $M=\sum_{p:(p,N^{(r)}_{p,\max})\in\mathcal{A}^{(r)}}N^{(r)}_{p,\max}$ is the total cost of sampling all such maximal arms once. This policy maintains uniform coverage over prompts while respecting cost asymmetries and ensures that the maximum scale of every prompt has an equal number of samples.
We now derive [^1] Proof is in the appendix. error bounds for Psst under the allocation policies described above.
@@
Let $R=\lceil\log_{2}|\mathcal{A}|\rceil$ be the number of trimming rounds, assume $o_{k}^{\min},o_{k}^{\max}]=[-1,1]$, and define the cost–gap complexity
$H_{1}^{c}=\max_{(c,a^{c,i})\neq(c,a^{c,1})}\frac{\bar{N}_{\max}}{\Delta_{c,a^{c,i}}^{\,2}},\quad H_{1}=\max_{c}H_{1}^{c}.$
$H_{2}^{c}=\max_{(c,a^{c,i})\neq(c,a^{c,1})}\frac{i\bar{N}_{\max}}{\Delta_{c,a^{c,i}}^{\,2}},\quad H_{2}=\max_{c}H_{2}^{c}.$
where $\Delta_{c,a^{c,i}}=Q^{\alpha}(c,a^{c,1})-Q^{\alpha}(c,a^{c,i})$. Under a context $c$, arms are indexed based on ascending order of $Q^{\alpha}(c,a)$ and $\bar{N}_{\max}$ is defined as $\frac{N(a^{c,1})+N_{\max}}{2}$. Here, $N(a^{c,i})$ is the number of completions generated by the i-th indexed arm. Running Psst with the structure-aware allocation for a total completion budget $T$ returns the optimal arm in every context with probability at least
$1-3|\mathcal{C}|R\exp\!\Big(-\tfrac{T}{\min(2|\mathcal{P}|H_{1},8|\mathcal{C}|H_{2})R}\Big).$
@@
$T=O\Big(\min(|\mathcal{P}|H_{1},|\mathcal{C}|H_{2})R\,\log\!\Bigl(\tfrac{|\mathcal{C}|R}{\delta}\Bigr)\Big).$
Note that applying Sequential-Halving without leveraging the structure of Iapo—specifically, without any form of information sharing across scales or contexts—incurs a sample complexity larger by a factor of $O(|\mathcal{C}|N_{\max})$.
@@
While we describe the algorithm assuming that each round uses a fresh dataset $\mathcal{D}$, it has been shown in similar halving-style algorithms (Fabiano and Cazenave 2021) that aggregating observations from all previous rounds—known as stockpiling—can improve the complexity of $T$ by reducing the outer $R$-factor, and we recommend using it with Psst.
@@
To further reduce the budget requirement of Psst, we introduce Top-$\boldsymbol{K}$ screening, a practical heuristic that executes a short, uniform prompt screening at unit scale to trim clearly suboptimal prompts before running full Psst. Top-$\boldsymbol{K}$ screening takes a budget fraction $T_{0}=\lfloor\rho T\rfloor$ ($\rho\in(0,1)$) from Psst. With scale restriction of N=1N{=}1, the budget is allocated uniformly across prompts: each $p\in\mathcal{P}$ receives $\big\lfloor T_{0}/|\mathcal{P}|\big\rfloor$ i.i.d. samples. Based on this data, $Q^{\alpha}(c,p,1)$ is estimated $\forall c\in\mathcal{C},p\in\mathcal{P}$. For each context $c$, we retain the $K$ best prompts $\mathcal{P}^{(0)}_{c}=\mathrm{Top}\text{-}K\{\,\widehat{Q}^{\alpha}(c,p,1):p\in\mathcal{P}\,\}$ and discard the rest. The subsequent Psst run is then restricted to the reduced arm sets $\mathcal{A}^{(1)}_{c}=\{(p,N):p\in\mathcal{P}^{(0)}_{c},\ N\in[N_{\max}]\}$ for each $c$, and uses the remaining budget $T^{\prime}=T-T_{0}$. In the next section, we demonstrate that the screening strategy can significantly improve performance in low training budget settings without compromising quality for practical tasks. However, theoretical guarantees comparable to those of full Psst cannot be established; counterexample tasks can be carefully constructed within the Iapo framework, where Top-$\boldsymbol{K}$ screening will behave suboptimally for any $K<|\mathcal{P}|$ (see results from synthetic environments).
@@
  • To demonstrate that Psst and the Top-$\boldsymbol{K}$ screening heuristic are highly effective at learning the policy $\pi_{T}$.
  • To show that Iapo improves the average cost-adjusted reward (ACR) on test queries compared to inference strategy agnostic optimization.
@@
We compare Psst and Top-$\boldsymbol{K}$ screening with several baselines. We denote Top-$\boldsymbol{K}$ screening with $K\in\{1,4,8\}$ as Psst+K1+K1, Psst+K4+K4, and Psst+K8+K8 respectively. For these heuristics, we fix $\rho=0.2$, which was found to perform best across all datasets using a sweep over $\rho\in\{0.05,0.1,0.2,0.3,0.4\}$. Full Psst is parameter-free and does not require any tuning. In our first set of experiments, we compare our proposed methods against several standard exploration strategies:
@@
  • UCB: At each turn, selects the arm with the highest optimistic $Q$ estimate. The exploration constant is $0.1$ after tuning.
@@
Note that all hyperparameter sweep results are in the appendix; we report results with the best setting found across all six datasets.
@@
Environments $\alpha$ $\lvert\mathcal{P}\rvert$ $N_{\max}$ $o_{k}^{\max}$ $|\mathcal{X}|$ $|\mathcal{C}|$
@@
We evaluated inference-aware optimization across a total of six environments. Key details are provided in Table 1. Environments 1 and 4 are synthetically constructed to mimic Iapo tasks, where prompt-query pair score distributions $O_{k}(x,p,\cdot)$ are modeled using categorical distributions. We introduce them to validate some of the theoretical findings. The remaining four environments are based on widely-used real-world datasets. Among these, MATH (Hendrycks et al. 2021) and CommonsenseQA (Talmor et al. 2018) are used to evaluate reasoning tasks under Majority Voting (MV), while Helpful-Harmless (Bai et al. 2022) and Summarization (Stiennon et al. 2020) are chosen for Best-of-N (BoN) evaluation.
For the MV tasks, the task objective is defined as an exact match with the correct answer. All three BoN tasks are bi-objective, and we use publicly available reward models from previous multi-objective LLM alignment studies to score completions (see appendix for links). The cost objective in all six tasks is defined to be proportional to the average number of tokens per response. For context specification, MV tasks include a budget regime $\{low,mid,high\}$, while BoN tasks include both the budget and the bi-objective weights, which range from 0.1 to 0.9 for each objective. For example, in the helpful-harmless task, a context might be represented as $\{\text{helpful}:0.3,\text{harmless}:0.7,\text{budget}:\text{high}~(1.0)\}$. Finally, for all environments, we set $N_{\max}$ to 32 because utility improvement diminishes sharply beyond $N=16$ across benchmarks for both BoN and MV.
To construct the environments, we first generated a set of instruction prompts for each task using ChatGPT-o3. We then generated 128 responses for each prompt–query pair and estimated the score distribution using a categorical model. All completions were produced using the Llama-3.3-70B-Instruct, a widely used open-source model (Grattafiori et al. 2024), which we treat as a black-box throughout our experiments. Generation was carried out with vLLM (Kwon et al. 2023) on a cluster of 8 A100 GPUs, totaling approximately 2,000 GPU-hours. Once the environments are constructed, all experiments can be run quickly via a standard CPU. We will publish the environments and code with the paper, enabling full reproducibility without any substantial computational resources.
@@
All reported curves are averages over 200 independent runs. For synthetic environments, we instantiate 200 independent environments and report the average performance across them. For the remaining four environments, each run reshuffles the dataset, performs an $80/20$ train–test split, and trains the policy on the training set. In all six environments, we evaluate ACR on the test set using $10,000$ samples. Performance for each budget is the mean across the 200 runs, with standard error of the mean (SEM) error bars. Statistical significance is assessed using the Wilcoxon paired two-sided test (Wilcoxon1945IndividualCB) with alpha = 0.05, and we indicate when differences are significant in the discussion. The full set of results is in the appendix.
Comparison of Exploration Strategies (Figure 4).
@@
Psst and the Top-$\boldsymbol{K}$ screening heuristic consistently outperform all baselines. Across all six domains, where the per-context action spaces are large ($|\mathcal{P}|N_{\max}\in[640,1536]$), UCB, softmax, and $\varepsilon$-greedy methods struggle to explore effectively. Among the baselines, UCB performs comparably in some domains after $T=20\text{K}$, but only with extensive hyperparameter tuning. Furthermore, these baselines are fully sequential and cannot leverage the cost and computational efficiency benefits of batch exploration. Full Psst attains the best final performance across four settings, while Top-$\boldsymbol{K}$ screening typically reaches strong policies faster, matching or exceeding Psst on three of the four real‑data tasks when the budget is small. Under aggressive pruning (small $K$), however, the heuristic becomes suboptimal—most notably on summarization and on the synthetic benchmarks—suggesting that Top-$\boldsymbol{K}$ screening is attractive under tight budgets, whereas full Psst is preferable for critical tasks such as long‑horizon, high‑frequency deployment. Finally, the statistical test also indicates that Psst, along with Top-$\boldsymbol{K}$ screening, significantly outperforms baselines in all six datasets and under nearly all budgets. These findings indicate that our approach reliably discovers well-aligned solutions using as few as $5\text{K}$ inference calls in practical settings.
Importance of Inference-Awareness (Figure 5).
@@
We examine the role of inference awareness in prompt optimization. Across all six datasets, Iapo methods markedly outperform the inference-agnostic methods, demonstrating the gains achievable when jointly optimizing the prompt and inference scale. TRIPLE (N= 1) fails as it does not leverage inference scaling. On the other hand, TRIPLE (N= Random) fails because it does not optimize the scaling for different contexts. The screening variant Psst+K1+K1—which effectively approximates a near‑decoupled (prompt‑only) procedure—fails to reach the optimum in most cases, performing competitively only on CommonsenseQA and showing pronounced underperformance on summarization. This is because it gets stuck with deceptive prompts that fail to scale compared to prompts that may not perform well under single-shot but improve significantly under scaling. These findings underscore the essential role of Iapo in aligning black‑box LLMs and the pitfalls of disjoint optimization. Overall, Iapo outperforms disjoint optimization by up to 25% and prompt-only optimization by up to 50% in our experiments.
@@
We present an inference-aware prompt optimization (Iapo) framework for aligning black-box LLMs, emphasizing that prompt design and deployment‑time inference scaling strategies are tightly coupled and should be optimized jointly. Our proposed Psst and Top-$\boldsymbol{K}$ screening heuristic demonstrate consistent improvements over strong baselines across six different settings. Looking ahead, we plan to explore richer inference scaling policies (e.g., tree search and parallel thinking). We also aim to extend the framework to multi-objective alignment with hard latency constraints and to study long-horizon deployments under distribution shift.
@@ Acknowledgments
Let $R=\lceil\log_{2}|\mathcal{A}|\rceil$ be the number of trimming rounds, assume $o_{k}^{\min},o_{k}^{\max}]=[-1,1]$, and define the cost–gap complexity
$H_{1}^{c}=\max_{(c,a^{c,i})\neq(c,a^{c,1})}\frac{\bar{N}_{\max}}{\Delta_{c,a^{c,i}}^{\,2}},\quad H_{1}=\max_{c}H_{1}^{c},$
$H_{2}^{c}=\max_{(c,a^{c,i})\neq(c,a^{c,1})}\frac{i\bar{N}_{\max}}{\Delta_{c,a^{c,i}}^{\,2}},\quad H_{2}=\max_{c}H_{2}^{c},$
where $\Delta_{c,a^{c,i}}=Q^{\alpha}(c,a^{c,1})-Q^{\alpha}(c,a^{c,i})$. Under a context $c$, arms are indexed based on ascending order of $Q^{\alpha}(c,a)$ and $\bar{N}_{\max}$ is defined as $\frac{N(a^{c,1})+N_{\max}}{2}$. Here, $N(a^{c,i})$ is the number of completions generated by the $i$-th indexed arm. Running Psst with the structure-aware allocation for a total completion budget $T$ returns the optimal arm in every context with probability at least
$1-3|\mathcal{C}|R\exp\!\left(-\frac{T}{\min\!\bigl(2|\mathcal{P}|H_{1},\;8|\mathcal{C}|H_{2}\bigr)\,R}\right).$
@@
$T=O\!\left(\min\!\bigl(|\mathcal{P}|H_{1},\;|\mathcal{C}|H_{2}\bigr)\,R\,\log\!\Bigl(\tfrac{|\mathcal{C}|R}{\delta}\Bigr)\right).$
Indexing convention.
Although Theorem 2 states that arms are indexed in ascending order of $Q^{\alpha}(c,a)$, throughout the proofs below we adopt the equivalent convention that $a^{c,1}$ denotes an optimal arm for context $c$, i.e., $Q^{\alpha}(c,a^{c,1})=\max_{a}Q^{\alpha}(c,a)$, and thus $Q^{\alpha}(c,a^{c,1})\geq Q^{\alpha}(c,a^{c,2})\geq\cdots$. All quantities (e.g., $\Delta_{c,a^{c,i}}$ and $\bar{N}_{i}$) are interpreted under this convention. In particular, $\Delta_{c,a^{c,i}}\geq 0$ for all $i\geq 2$.
@@
Fix a context $c$ and recall the above convention that $a^{c,1}$ denotes an optimal arm for $c$. Assume $a^{c,1}$ was not eliminated before round $r$. Let $t_{r,i}$ denote the number of i.i.d. samples used to estimate $Q^{\alpha}(c,a^{c,i})$ in round $r$, and define the (effective) harmonic mean
$\operatorname{harmonic}(u,v)\;:=\;\left(u^{-1}+v^{-1}\right)^{-1}.$
Then, by Hoeffding’s inequality, for any arm $a^{c,i}\in\mathcal{A}_{c}^{(r)}$,
$\Pr\!\big[\hat{Q}^{\alpha,(r)}(c,a^{c,1})$ $<\hat{Q}^{\alpha,(r)}(c,a^{c,i})\big]$
$\leq\exp\!\Bigl(-\tfrac{1}{2}\,\operatorname{harmonic}(t_{r,1},t_{r,i})\,\Delta_{c,a^{c,i}}^{2}\Bigr).$
Let $N_{r}$ be the number of arms in $\mathcal{A}_{c}^{(r)}$ whose empirical estimate exceeds that of $a^{c,1}$. Then
$\mathbb{E}[N_{r}]$ $\leq\sum_{a^{c,i}\in\mathcal{A}_{c}^{(r)}}\Pr\!\big[\hat{Q}^{\alpha,(r)}(c,a^{c,1})<\hat{Q}^{\alpha,(r)}(c,a^{c,i})\big]$
$\leq\sum_{a^{c,i}\in\mathcal{A}_{c}^{(r)}}\exp\!\Bigl(-\tfrac{1}{2}\,\operatorname{harmonic}(t_{r,1},t_{r,i})\,\Delta_{c,a^{c,i}}^{2}\Bigr)$
$\leq\sum_{a^{c,i}\in\mathcal{A}_{c}^{(r)}}\exp\!\Bigl(-\Delta_{c,a^{c,i}}^{2}\cdot\tfrac{T}{2|\mathcal{P}|\,\bar{N}_{i}\,\log_{2}|\mathcal{A}|}\Bigr)$
$\leq|\mathcal{A}_{c}^{(r)}|\,\max_{a^{c,i}\in\mathcal{A}_{c}^{(r)}}\exp\!\Bigl(-\Delta_{c,a^{c,i}}^{2}\cdot\tfrac{T}{2|\mathcal{P}|\,\bar{N}_{\max}\,\log_{2}|\mathcal{A}|}\Bigr)$
$\leq|\mathcal{A}_{c}^{(r)}|\,\exp\!\Bigl(-\tfrac{T}{2|\mathcal{P}|\,H_{1}^{c}\,R}\Bigr),$
where $\bar{N}_{i}:=\frac{N(a^{c,1})+N(a^{c,i})}{2}\leq\bar{N}_{\max}$ and we used $R=\lceil\log_{2}|\mathcal{A}|\rceil$.
For the best arm to be eliminated in round $r$, at least half of the active arms must (incorrectly) rank above it, i.e., $N_{r}\geq\tfrac{1}{2}\,|\mathcal{A}_{c}^{(r)}|$. By Markov’s inequality,
$\Pr\!\Big[N_{r}\geq\tfrac{1}{2}|\mathcal{A}_{c}^{(r)}|\Big]\;\leq\;\frac{2\,\mathbb{E}[N_{r}]}{|\mathcal{A}_{c}^{(r)}|}\;\leq\;2\,\exp\!\Bigl(-\tfrac{T}{2|\mathcal{P}|H_{1}^{c}R}\Bigr),$
which proves the lemma. ∎